OpenStudy (anonymous):

Solve x^2 - 8x + 15 = 0

OpenStudy (anonymous):

can you think of two numbers that multiply to be 15 and add to be -8?

OpenStudy (anonymous):

3 and 5

OpenStudy (anonymous):

so 3 and 5 add up to be -8?

hero (hero):

x^2 - 3x - 5x + 15 = 0 x(x-3) -5(x-3) = 0 (x-3)(x-5) = 0